OpenStudy (anonymous):

Can someone please show me how to work this out? For each of the following quadratic relationships: Find the transformation of the graph y=x2 (two meant to be square) that have taken place So, I don't understand the meaning of transformation??

OpenStudy (anonymous):

it just means to tell how it changed from the original parent funtion

OpenStudy (anonymous):

function*

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

for example: y=x^2 + 2 would be an upward shift of two units

OpenStudy (anonymous):

oh yes. What about something like y=(2(x+1)+3

OpenStudy (anonymous):

Im mean y=2(x+1)+3

OpenStudy (anonymous):

what is the original?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

is it supposed to be \[y=(x+1)^{2}+3\]

OpenStudy (anonymous):

oh sorry it is supposed to be y=2(x-1)2+3

OpenStudy (anonymous):

alright. What does the plus three do?

OpenStudy (anonymous):

it goes up?

OpenStudy (anonymous):

how many units?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

good. how about the -1 inside the parenthesis

OpenStudy (anonymous):

it goes down by 1 unit?

OpenStudy (anonymous):

no, if it is going to go down, it would be outside the parenthesis like the three

OpenStudy (anonymous):

or does it go to the right 5 units ?

OpenStudy (anonymous):

sorry I meant right 1 unit?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and what will the two do?

OpenStudy (anonymous):

I don't know what it does now

OpenStudy (anonymous):

do you multiply them?

OpenStudy (anonymous):

You do, but what happens when you multiply x by a number?

OpenStudy (anonymous):

we know what happens when we add a number outside parentheses and when we add one inside, how about multiplication?

OpenStudy (anonymous):

well if say you are multiplying x by 48 then it would be 48x?

OpenStudy (anonymous):

yes, but what happens to the graph of a parabola when you multiply the x^2 term

OpenStudy (anonymous):

I don't know

OpenStudy (anonymous):

it will make it taller and skinnier, or shorter and fatter

OpenStudy (anonymous):

http://www.purplemath.com/modules/fcntrans2.htm look at this page- it might help some!

OpenStudy (anonymous):

does it make it shorter and fatter?

OpenStudy (anonymous):

okay thank you I will have a look at that. I need to go now but I will be back online

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

It will be taller and skinnier btw